OpenStudy (anonymous):

find the differential of f(x)=x^(1/2) with respect to x and dx

OpenStudy (ajprincess):

\(\Large f(x)=x^n\) \(\Large f'(x)=nx^{(n-1)}dx\) Does that help? @motlatsi

OpenStudy (anonymous):

i dont get it man, may you help out

OpenStudy (ajprincess):

ya. if f(x) is of the form x^n. then its differentiation will be in the form f'(x)=nx^(n-1)dx. Let me give an example. suppose f(x)=x^3 so n=3 here. f'(x)=3x^(3-1)dx =3x^2dx. Getting it? @motlatsi